OpenStudy (anonymous):

The position vector of the point A and B are 2i+3j and -i+5j.find the position of vector of the point C such that AC=2AB. Calculate the angle between vector AB and OB.

OpenStudy (ajprincess):

OA=2i+3j and OB= -i+5j AB=OB-OA =-i+5j-2i-3j =-3i+2j AC=OC-OA =2AB 2AB=OC-OA 2(-3i+2j)=OC-2i-3j -6i+4j+2i+3j=OC  Use this to find the angle.